OpenStudy (anonymous):

Solutions to this equation: 4x+5y=20 x+y=1

OpenStudy (anonymous):

Multiply the second equation by either 4 or 5 and subtract it from the first equation to eliminate one of the variables x or y , then solve.... so if I multiply Eq2 by 4 it becomes 4x+4y=4 subtract this from Eq1 to eliminate x gives 4x-4x+5y-4y=20-4 y=16 so to solve for x substitue y =16 in Eq2 x + 16=1; therefore x=-15